Find the equation of the line through the given pair of points in standard form using only integers. (4,-1) and (-2,-6)
Answer by checkley77(12844)

(4,-1) and (-2,-6)
First we find the slope (m)
(y2-y1)/(x2-x1)
(-6+1)/(-2-4)=-5/-6=5/6
Y=mX+b
-1=4*5/6+b
-1=20/6+b
b=-1-20/6
b=(-6-20)/6
b=-26/6
y=5x/6-26/6
6y=5x-26 ans. using integers.